# SynPO: Synergizing Descriptiveness and Preference Optimization for Video Detailed Captioning

### v1

@@ 5.2.1 Comparison with DPO and Various Variants
SynPO-v1 $-\sigma\left(\alpha\cdot\exp\left(\overline{\log S(y_{w})}\right)-\alpha\cdot\exp\left(\overline{\log S(y_{l})}\right)\right)$
SynPO-v2 $-\sigma\left(\alpha\cdot\exp\left(\overline{\log S(y_{w})}\right)-\alpha\cdot\exp\left(\overline{\log S(y_{l})}\right)\right)-\beta\,\overline{\log S(y_{w})}$
SynPO-v3 $-\sigma\left(\alpha\cdot\overline{S(y_{w})}-\alpha\cdot\overline{S(y_{l})}\right)-\beta\,\overline{S(y_{w})}$
SynPO-v4 $-\sigma\left(\alpha\cdot\overline{S(y_{w})}-\alpha\cdot\overline{S(y_{l})}\right)-\beta\,\overline{\log S(y_{w})}$
SynPO-v5 $-\left(\alpha\cdot\exp\left(\overline{\log S(y_{w})}\right)-\alpha\cdot\exp\left(\overline{\log S(y_{l})}\right)\right)-\beta\,\overline{S(y_{w})}$
@@ 5.2.2 Effectiveness in NLP Domain
Training Recipe. Training experiments are conducted using Llama3-8B [llama3modelcard] (Base and Instruct) and Mistral-7B [Jiang2023Mistral7] (Base and Instruct). For both Llama3-8B-Base and Mistral-7B-Base, we employ a training pipeline [tunstall2023zephyr]. First, we train a base model on the UltraChat-200k dataset [Ding2023EnhancingCL] to obtain an SFT model. Then, we perform preference optimization on the UltraFeedback dataset [Cui2024UltraFeedbackBL] using the SFT model as the starting point. For Llama3-8B-Instruct and Mistral-7B-Instruct, we implement an on-policy evaluation strategy following SimPO [meng2024simposimplepreferenceoptimization]. Specifically, prompts from UltraFeedback are used to regenerate positive and negative preference pairs via SFT models. For each prompt, five responses are sampled from the SFT model and rank them using PairRM (LLM-Blender) [jiang2023llm]. The highest-ranked response is selected as the positive preference, and the lowest-ranked response is designated as the negative one.
Evaluation Benchmark. Building upon recent methodologies in preference-based fine-tuning [rafailov2024direct, tunstall2023zephyr], we evaluate model performance using standardized frameworks. These include both versions of the HuggingFace Open LLM Leaderboard [eval-harness], summarized in Table 6, and comprehensive instruction-following benchmarks (AlpacaEval2 and MT-Bench) as reported in Table 5. Detailed descriptions of evaluation tasks and procedures are provided in Appendix A.
Table 5: AlpacaEval2 [AlpacaEval] and MT-Bench [zheng2023judging] results under the four settings. LC and WR denote length-controlled and raw win rate, respectively.
@@
DPO [rafailov2024direct] 15.1 12.5 7.3 26.8 24.9 7.6 18.2 15.5 7.7 40.3 37.9 8.0
DPOP [pal2024smaugfixingfailuremodes] 16.1 12.8 7.4 27.1 24.6 7.7 16.7 14.3 7.6 45.2 39.1 8.2
IPO [Azar2023AGT] 11.8 9.4 7.2 20.3 20.3 7.8 14.4 14.2 7.4 35.6 35.6 8.3
KTO [Ethayarajh2024KTOMA] 13.1 9.1 7.0 24.5 23.6 7.7 14.2 12.4 7.8 33.1 31.8 8.2
CPO [xu2024contrastive] 9.8 8.9 6.8 23.8 28.8 7.5 10.8 8.1 7.4 28.9 32.2 8.0
SimPO [meng2024simposimplepreferenceoptimization] 21.5 20.8 7.3 32.1 34.8 7.6 22.0 20.3 7.7 44.7 40.5 8.0
@@
Table 6: Evaluation results on various tasks from the Huggingface Open Leaderboards [open-llm-leaderboard-v1, open-llm-leaderboard-v2] show that our SynPO achieves superior or comparable performance to other.
@@ Appendix A The Evaluation Benchmarks
MMLU-PRO. [wang2024mmluprorobustchallengingmultitask] It is a robust benchmark for evaluating cross-disciplinary reasoning in LLMs, comprising 12,000 complex questions spanning STEM, humanities, and professional domains, with ten answer options per question to minimize random guessing and emphasize analytical depth. It integrates problems from diverse sources (e.g., original MMLU, TheoremQA, SciBench), employs chain-of-thought reasoning requirements, and demonstrates enhanced robustness to prompt variations, as evidenced by leading models achieving  71% accuracy while highlighting significant performance gaps compared to earlier benchmarks.
@@
IFEval. [zhou2023instructionfollowingevaluationlargelanguage] It is a benchmark dataset designed to evaluate the in-context learning and few-shot reasoning capabilities of LLMs across diverse NLP tasks, featuring carefully curated prompts and annotations to assess performance under varying input conditions and task complexities.
@@
BBH. [suzgun2022challengingbigbenchtaskschainofthought] Big Bench Hard is a benchmark dataset designed to evaluate the cross-domain reasoning capabilities of LLMs, comprising 23 high-difficulty tasks that emphasize multi-step logical deduction, attention control, and memory retention, with a focus on few-shot learning scenarios and the application of chain-of-thought (CoT) reasoning to challenge models beyond their standard performance thresholds.
@@
HellaSwag. [zellers2019hellaswagmachinereallyfinish] It is a NLP benchmark designed to evaluate machine commonsense reasoning and contextual understanding, featuring more than 100,000 context-rich question-answer pairs generated via crowdsourcing and adversarial filtering, with a focus on challenging models to infer plausible continuations of text beyond superficial pattern matching.
@@
WinoGrande. [sakaguchi2019winograndeadversarialwinogradschema] It is a benchmark dataset designed to evaluate the commonsense reasoning and pronoun disambiguation capabilities of LLMs, extending the Winograd Schema Challenge by introducing 44,000 context-dependent questions with multiple-choice answers that require resolving ambiguous references through deep contextual understanding and implicit world knowledge.
@@
TruthfulQA. [lin-etal-2022-truthfulqa] It is designed to evaluate the factual accuracy and truthfulness of LLMs, comprising 817 adversarially crafted zero-shot questions across 38 topics with verified true/false answers, emphasizing the model’s ability to avoid generating false statements through rigorous human-validated sources and challenging high-probability training-distribution biases.
@@
GSM8K. [cobbe2021trainingverifierssolvemath] It is a benchmark dataset designed to evaluate the multistep arithmetic reasoning capabilities of NLP models, comprising 8,500 high-quality grade-school-level math word problems that require 2–8 sequential operations using basic arithmetic, with answers presented in annotated natural language formats to facilitate both model training and rigorous assessment of mathematical problem-solving robustness.
@@
ARC-C. [clark2018thinksolvedquestionanswering] The ARC dataset consists of 7,787 science questions, all non-diagram, multiple choice (tpically 4-way multiple choice). They are drawn from a variety of sources, and sorted into a challenge set of 2,590 “hard" questions (those that both a retrieval and a co-occurrence method fail to answer correctly) and an easy set of 5,197 questions. Questions vary in their target student grade level (as assigned by the examiners who authored the questions), ranging from 3rd grade to 9th.
@@ Appendix B Details of DPO Variants
DPO $-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$ $\beta\in\{0.01,0.05,0.1\}$
@@
CPO $-\log\sigma\left(\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)\right)-\lambda\log\pi_{\theta}(y_{w}|x)$ $\lambda=1.0,\,\,\beta\in\{0.01,0.05,0.1\}$
DPOP $-\left[\log\sigma\left(\beta\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\lambda\cdot\max\left(0,\log\frac{\pi_{\text{ref}}(y_{w}|x)}{\pi_{\theta}(y_{w}|x)}\right)\right)\right)\right]$ $\beta\in\{0.5,0.1,0.2,0.3\}$, $\lambda\in\{5,10,25,50\}$
KTO $&-\lambda_{w}\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-z_{\text{ref}}\right)+\lambda_{l}\sigma\left(z_{\text{ref}}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right),\\ &\text{where }z_{\text{ref}}=\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\beta\text{KL}\left(\pi_{\theta}(y|x)\|\pi_{\text{ref}}(y|x)\right)\right]\end{aligned}$ $\lambda_{l}=\lambda_{w}=1.0$, $\beta\in\{0.01,0.05,0.1\}$
SimPO $-\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)$ $\beta\in\{2.0,2.5\}$
@@
SynPO $-\left[\sigma\left(\alpha\cdot\exp\left(\,\overline{\log S(y_{w})}\right)-\alpha\cdot\exp\left(\,\overline{\log S(y_{l})\,}\right)\right)+\beta\cdot\overline{S(y_{w})}\right]$ $\alpha\in\{20,30,50\}$, $\beta\in\{0.1,0.2,0.3\}$
DPO. Direct Preference Optimization [rafailov2024direct] uses log-likelihood differences to implicitly represent the reward function, eliminating the need for explicit reward model like RLHF. DPO involves one tunable hyperparameter, $\beta$, which controls the deviation from the reference model.
IPO. Identity Preference Optimization [azar2024general] minimizes a squared loss regression problem by defining an alternative reward function, avoiding unstable RL training. IPO involves one hyperparameter, $\beta$, to adjust the reward margin.
CPO. Contrastive Preference Optimization [xucontrastive] uses log-likelihood as the reward and is trained alongside a Supervised Fine-Tuning (SFT) objective. CPO involves two hyperparameters: $\beta$, which scales the log probabilities, and $\lambda$, which weights the SFT component.
SimPO. Simple Preference Optimization [meng2024simposimplepreferenceoptimization] eliminates the need for a reference model and optimizes a length-regularized probability of response pairs. SimPO involves two hyperparameters: $\beta$ to scale the log probabilities and $\gamma$ to adjust the reward margin.
KTO. Kahneman-Tversky Optimization [ethayarajh2024kto] learns from non-paired preference data. KTO involves three hyperparameters: $\beta$, which controls the deviation from the reference model; $\lambda_{w}$ and $\lambda_{l}$, which weight the preference components for winning and losing responses, respectively.
DPOP. DPO-Positive [pal2024smaugfixingfailuremodes] adds a new term to the loss which leads every token to be incentivised toward the preferred completion.
@@
SynPO-v1 $-\sigma\left(\alpha\cdot\exp\left(\overline{\log S(y_{w})}\right)-\alpha\cdot\exp\left(\overline{\log S(y_{l})}\right)\right)$ $\alpha\in\{20,30,50\}$
SynPO-v2 $-\sigma\left(\alpha\cdot\exp\left(\overline{\log S(y_{w})}\right)-\alpha\cdot\exp\left(\overline{\log S(y_{l})}\right)\right)-\beta\,\overline{\log S(y_{w})}$ $\alpha\in\{20,30,50\}$, $\beta\in\{0.05,0.1,0.2,0.3\}$
SynPO-v3 $-\sigma\left(\alpha\cdot\overline{S(y_{w})}-\alpha\cdot\overline{S(y_{l})}\right)-\beta\,\overline{S(y_{w})}$ $\alpha\in\{10,20,30,50\}$, $\beta\in\{0.1,0.2,0.3\}$
SynPO-v4 $-\sigma\left(\alpha\cdot\overline{S(y_{w})}-\alpha\cdot\overline{S(y_{l})}\right)-\beta\,\overline{\log S(y_{w})}$ $\alpha\in\{10,20,30,50\}$, $\beta\in\{0.05,0.1,0.2,0.3\}$
SynPO-v5 $-\left(\alpha\cdot\exp\left(\overline{\log S(y_{w})}\right)-\alpha\cdot\exp\left(\overline{\log S(y_{l})}\right)\right)-\beta\,\overline{S(y_{w})}$ $\alpha=1$, $\beta\in\{0.01,0.02,0.05,0.1,0.15,0.2\}$

### v2

@@ 5.2.1 Comparison with DPO and Various Variants
SynPO-v1 $\sigma\left(\alpha\cdot\exp\left(\overline{\log S(y_{w})}\right)-\alpha\cdot\exp\left(\overline{\log S(y_{l})}\right)\right)$
SynPO-v2 $\sigma\left(\alpha\cdot\exp\left(\overline{\log S(y_{w})}\right)-\alpha\cdot\exp\left(\overline{\log S(y_{l})}\right)\right)-\beta\,\overline{\log S(y_{w})}$
SynPO-v3 $\sigma\left(\alpha\cdot\overline{S(y_{w})}-\alpha\cdot\overline{S(y_{l})}\right)-\beta\,\overline{S(y_{w})}$
SynPO-v4 $\sigma\left(\alpha\cdot\overline{S(y_{w})}-\alpha\cdot\overline{S(y_{l})}\right)-\beta\,\overline{\log S(y_{w})}$
SynPO-v5 $\left(\alpha\cdot\exp\left(\overline{\log S(y_{w})}\right)-\alpha\cdot\exp\left(\overline{\log S(y_{l})}\right)\right)-\beta\,\overline{S(y_{w})}$
@@ 5.2.2 Effectiveness in NLP Domain
Training Recipe. Training experiments are conducted using Llama3-8B [undefb] (Base and Instruct) and Mistral-7B [undefad] (Base and Instruct). For both Llama3-8B-Base and Mistral-7B-Base, we employ a training pipeline [undefaak]. First, we train a base model on the UltraChat-200k dataset [undefu] to obtain an SFT model. Then, we perform preference optimization on the UltraFeedback dataset [undefr] using the SFT model as the starting point. For Llama3-8B-Instruct and Mistral-7B-Instruct, we implement an on-policy evaluation strategy following SimPO [undefax]. Specifically, prompts from UltraFeedback are used to regenerate positive and negative preference pairs via SFT models. For each prompt, five responses are sampled from the SFT model and rank them using PairRM (LLM-Blender) [undefae]. The highest-ranked response is selected as the positive preference, and the lowest-ranked response is designated as the negative one.
Evaluation Benchmark. Building upon recent methodologies in preference-based fine-tuning [undefaac, undefaak], we evaluate model performance using standardized frameworks. These include both versions of the HuggingFace Open LLM Leaderboard [undefaa], summarized in Table 6, and comprehensive instruction-following benchmarks (AlpacaEval2 and MT-Bench) as reported in Table 5. Detailed descriptions of evaluation tasks and procedures are provided in Appendix A.
Table 5: AlpacaEval2 [undefao] and MT-Bench [undefaaad] results under the four settings. LC and WR denote length-controlled and raw win rate, respectively.
@@
DPO [undefaac] 15.1 12.5 7.3 26.8 24.9 7.6 18.2 15.5 7.7 40.3 37.9 8.0
DPOP [undefaaa] 16.1 12.8 7.4 27.1 24.6 7.7 16.7 14.3 7.6 45.2 39.1 8.2
IPO [undefd] 11.8 9.4 7.2 20.3 20.3 7.8 14.4 14.2 7.4 35.6 35.6 8.3
KTO [undefv] 13.1 9.1 7.0 24.5 23.6 7.7 14.2 12.4 7.8 33.1 31.8 8.2
CPO [undefaax] 9.8 8.9 6.8 23.8 28.8 7.5 10.8 8.1 7.4 28.9 32.2 8.0
SimPO [undefax] 21.5 20.8 7.3 32.1 34.8 7.6 22.0 20.3 7.7 44.7 40.5 8.0
@@
Table 6: Evaluation results on various tasks from the Huggingface Open Leaderboards [undefh, undefy] show that our SynPO achieves superior or comparable performance to other.
@@ Appendix A The Evaluation Benchmarks
MMLU-PRO. [undefaar] It is a robust benchmark for evaluating cross-disciplinary reasoning in LLMs, comprising 12,000 complex questions spanning STEM, humanities, and professional domains, with ten answer options per question to minimize random guessing and emphasize analytical depth. It integrates problems from diverse sources (e.g., original MMLU, TheoremQA, SciBench), employs chain-of-thought reasoning requirements, and demonstrates enhanced robustness to prompt variations, as evidenced by leading models achieving  71% accuracy while highlighting significant performance gaps compared to earlier benchmarks.
@@
IFEval. [undefaaaf] It is a benchmark dataset designed to evaluate the in-context learning and few-shot reasoning capabilities of LLMs across diverse NLP tasks, featuring carefully curated prompts and annotations to assess performance under varying input conditions and task complexities.
@@
BBH. [undefaai] Big Bench Hard is a benchmark dataset designed to evaluate the cross-domain reasoning capabilities of LLMs, comprising 23 high-difficulty tasks that emphasize multi-step logical deduction, attention control, and memory retention, with a focus on few-shot learning scenarios and the application of chain-of-thought (CoT) reasoning to challenge models beyond their standard performance thresholds.
@@
HellaSwag. [undefaaaa] It is a NLP benchmark designed to evaluate machine commonsense reasoning and contextual understanding, featuring more than 100,000 context-rich question-answer pairs generated via crowdsourcing and adversarial filtering, with a focus on challenging models to infer plausible continuations of text beyond superficial pattern matching.
@@
WinoGrande. [undefaae] It is a benchmark dataset designed to evaluate the commonsense reasoning and pronoun disambiguation capabilities of LLMs, extending the Winograd Schema Challenge by introducing 44,000 context-dependent questions with multiple-choice answers that require resolving ambiguous references through deep contextual understanding and implicit world knowledge.
@@
TruthfulQA. [undefas] It is designed to evaluate the factual accuracy and truthfulness of LLMs, comprising 817 adversarially crafted zero-shot questions across 38 topics with verified true/false answers, emphasizing the model’s ability to avoid generating false statements through rigorous human-validated sources and challenging high-probability training-distribution biases.
@@
GSM8K. [undefq] It is a benchmark dataset designed to evaluate the multistep arithmetic reasoning capabilities of NLP models, comprising 8,500 high-quality grade-school-level math word problems that require 2–8 sequential operations using basic arithmetic, with answers presented in annotated natural language formats to facilitate both model training and rigorous assessment of mathematical problem-solving robustness.
@@
ARC-C. [undefp] The ARC dataset consists of 7,787 science questions, all non-diagram, multiple choice (tpically 4-way multiple choice). They are drawn from a variety of sources, and sorted into a challenge set of 2,590 “hard" questions (those that both a retrieval and a co-occurrence method fail to answer correctly) and an easy set of 5,197 questions. Questions vary in their target student grade level (as assigned by the examiners who authored the questions), ranging from 3rd grade to 9th.
@@ Appendix B Details of DPO Variants
DPO $\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$ $\beta\in\{0.01,0.05,0.1\}$
@@
CPO $\log\sigma\left(\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)\right)-\lambda\log\pi_{\theta}(y_{w}|x)$ $\lambda=1.0,\,\,\beta\in\{0.01,0.05,0.1\}$
DPOP $\left[\log\sigma\left(\beta\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\lambda\cdot\max\left(0,\log\frac{\pi_{\text{ref}}(y_{w}|x)}{\pi_{\theta}(y_{w}|x)}\right)\right)\right)\right]$ $\beta\in\{0.5,0.1,0.2,0.3\}$, $\lambda\in\{5,10,25,50\}$
KTO $\begin{aligned} &-\lambda_{w}\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-z_{\text{ref}}\right)+\lambda_{l}\sigma\left(z_{\text{ref}}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right),\\ &\text{where }z_{\text{ref}}=\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\beta\text{KL}\left(\pi_{\theta}(y|x)\|\pi_{\text{ref}}(y|x)\right)\right]\end{aligned}$ $\lambda_{l}=\lambda_{w}=1.0$, $\beta\in\{0.01,0.05,0.1\}$
SimPO $\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)$ $\beta\in\{2.0,2.5\}$
@@
SynPO $\left[\sigma\left(\alpha\cdot\exp\left(\,\overline{\log S(y_{w})}\right)-\alpha\cdot\exp\left(\,\overline{\log S(y_{l})\,}\right)\right)+\beta\cdot\overline{S(y_{w})}\right]$ $\alpha\in\{20,30,50\}$, $\beta\in\{0.1,0.2,0.3\}$
DPO. Direct Preference Optimization [undefaac] uses log-likelihood differences to implicitly represent the reward function, eliminating the need for explicit reward model like RLHF. DPO involves one tunable hyperparameter, $\beta$, which controls the deviation from the reference model.
IPO. Identity Preference Optimization [undefe] minimizes a squared loss regression problem by defining an alternative reward function, avoiding unstable RL training. IPO involves one hyperparameter, $\beta$, to adjust the reward margin.
CPO. Contrastive Preference Optimization [undefaaw] uses log-likelihood as the reward and is trained alongside a Supervised Fine-Tuning (SFT) objective. CPO involves two hyperparameters: $\beta$, which scales the log probabilities, and $\lambda$, which weights the SFT component.
SimPO. Simple Preference Optimization [undefax] eliminates the need for a reference model and optimizes a length-regularized probability of response pairs. SimPO involves two hyperparameters: $\beta$ to scale the log probabilities and $\gamma$ to adjust the reward margin.
KTO. Kahneman-Tversky Optimization [undefw] learns from non-paired preference data. KTO involves three hyperparameters: $\beta$, which controls the deviation from the reference model; $\lambda_{w}$ and $\lambda_{l}$, which weight the preference components for winning and losing responses, respectively.
DPOP. DPO-Positive [undefaaa] adds a new term to the loss which leads every token to be incentivised toward the preferred completion.
@@
SynPO-v1 $\sigma\left(\alpha\cdot\exp\left(\overline{\log S(y_{w})}\right)-\alpha\cdot\exp\left(\overline{\log S(y_{l})}\right)\right)$ $\alpha\in\{20,30,50\}$
SynPO-v2 $\sigma\left(\alpha\cdot\exp\left(\overline{\log S(y_{w})}\right)-\alpha\cdot\exp\left(\overline{\log S(y_{l})}\right)\right)-\beta\,\overline{\log S(y_{w})}$ $\alpha\in\{20,30,50\}$, $\beta\in\{0.05,0.1,0.2,0.3\}$
SynPO-v3 $\sigma\left(\alpha\cdot\overline{S(y_{w})}-\alpha\cdot\overline{S(y_{l})}\right)-\beta\,\overline{S(y_{w})}$ $\alpha\in\{10,20,30,50\}$, $\beta\in\{0.1,0.2,0.3\}$
SynPO-v4 $\sigma\left(\alpha\cdot\overline{S(y_{w})}-\alpha\cdot\overline{S(y_{l})}\right)-\beta\,\overline{\log S(y_{w})}$ $\alpha\in\{10,20,30,50\}$, $\beta\in\{0.05,0.1,0.2,0.3\}$
SynPO-v5 $\left(\alpha\cdot\exp\left(\overline{\log S(y_{w})}\right)-\alpha\cdot\exp\left(\overline{\log S(y_{l})}\right)\right)-\beta\,\overline{S(y_{w})}$ $\alpha=1$, $\beta\in\{0.01,0.02,0.05,0.1,0.15,0.2\}$